# Embodied Agents Meet Personalization: Exploring Memory Utilization for Personalized Assistance

### v1

@@ 4.3 Dataset Construction Process
Figure 3: The performance results without using episodic memory. Original indicates the conventional object rearrangement task episodes.
Step 4: Quality control.
To ensure data quality, we first heuristically filtered episodes containing similar memories referencing identical objects within scenes, preventing interference between similar episodic memories. Subsequently, we manually reviewed episodes from the memory acquisition stage where GPT-4o failed to successfully complete the task, where we filtered out episodes that contained unnatural instructions or cases where the generated instructions did not match the intended goal representation, ensuring the quality of our evaluation data.
4.4 Validation of Memento
To validate that Memento effectively assesses embodied agents’ memory utilization capability, we compare performance across stages in a setup without memory retrieval. As shown in Figure 3, compared to the results of the original object rearrangement task and the memory acquisition stage, embodied agents struggle to complete tasks in the memory utilization stage. Since the underlying episode remains the same across all tasks, this result confirms that the instruction $I_{util}$ in the memory utilization stage is difficult to interpret without access to previous interaction histories. Notably, we observed a particular behavior of the embodied agents: when there are two objects of the same type, agents that do not understand which one is the actual target tend to randomly select one and proceed with the task. This explains the reason why performance in the single-memory task is higher than that of the joint-memory task.
5 Evaluating Personalized Knowledge Utilization in Episodic Memory
5.1 Experimental Setup
Evaluation metrics.
Following Chang et al. [7], we use two main metrics: Percent Complete (PC) for the proportion of goal completion, and Success Rate (SR) for full task completion. We also report Sim Steps, which show the number of simulation steps required for agents to complete the task, and Planning Cycles, which indicate the number of LLM inference calls made during task execution. To evaluate memory utilization, we also report performance drops between acquisition and utilization stages as $\Delta$PC and $\Delta$SR. Note that for joint-memory tasks, these differences are computed relative to the average performance of the corresponding acquisition-stage episodes.
Implementations.
Following prior work [45, 38, 7], we implement a LLM-powered embodied agent architecture, where the LLM functions as a high-level policy planner that selects appropriate skills from a predefined skill library. We use ReAct [56] prompt format for LLMs to take actions. Additionally, we implement a top-5 memory retrieval setup  [^2] We use an embedding-based retrieval method with the all-mpnet-base-v2 Sentence Transformer [40]. for the memory utilization stage, ensuring the corresponding memory is included in the retrieved results by randomly replacing one memory if the correct one is not initially retrieved. Full implementation details are in AppendixB.2.
Models.
We evaluate embodied agents powered by a range of LLMs to compare memory utilization capabilities across model families and sizes, including proprietary models (GPT-4o [21], Claude-3.5-Sonnet [3]) and open-source models (Llama-3.1-70b/8b [15], Qwen-2.5-72b/7b [54]).
@@ 5.1 Experimental Setup
Model Stage Task Type (Memory) Planning Cycles $\downarrow$ Sim Steps $\downarrow$ Percent Complete $\uparrow$ $\Delta$ PC Success Rate $\uparrow$ $\Delta$ SR
@@
5.2 Main Results
LLM-powered embodied agents struggle with understanding personalized knowledge.
As shown in Table 1, while GPT-4o maintains a relatively high success rate in the single-memory task, all models show a success rate drop over 20% compared to the memory acquisition stage. In particular, for joint-memory tasks even GPT-4o exhibits a 30.5% drop in success rate, highlighting the increased difficulty of these settings. This substantial performance decline demonstrates that even frontier models struggle to accurately reference personalized knowledge from episodic memory, and often fail to consistently apply it across multiple steps in long-horizon task planning.
@@ 5.2 Main Results
LLM-powered embodied agents exhibit increased exploration behavior on joint-memory tasks.
Joint-memory task results reveal that LLMs (even GPT-4o) find it difficult to recall personalized knowledge from different memory sources. The number of planning cycles and simulation steps significantly increases compared to other tasks, this suggests that the embodied agent fails to correctly interpret the instruction, leading to excessive exploration during task execution. Also the performance gap between percent complete and success rate is larger than in the single-memory task, which indicates that the agent frequently misses part of the necessary information for successful task completion.
5.3 In-depth Analysis
To better understand the challenges of providing personalized assistance, we examine two key aspects: the root causes of their limitations and how the top-$k$ memory setting affects memory utilization.
5.3.1 Personalized Knowledge Type-based Analysis
We analyze the performance gap between the memory acquisition stage and the single-memory task from the memory utilization stage by comparing success rates across different types of personalized knowledge. Analysis results for the joint-memory task are provided in Appendix D.2.
LLMs can recall objects but struggle to comprehend action sequences.
As shown in Figure 4, all models exhibit substantially larger performance drops for tasks requiring understanding user patterns compared to object semantics. Notably, the performance drop in object semantics is minimal, indicating that LLMs are relatively effective at directly recalling relevant memories to identify target objects. In contrast, tasks involving user patterns pose a significantly greater challenge, as they require integrating and reasoning over sequences of events.
5.3.2 Memory Top-K Analysis
We analyze how varying the number of retrieved memories ($k$) impacts the agent’s ability to utilize personalized knowledge from episodic memory.
Irrelevant memories act as a distraction to agents.
In Figure 5, we observe that as $k$ increases, all models exhibit consistent performance degradation across both task types, highlighting the difficulty of identifying the exact information from a growing set of retrieved memories. Consistent with previous findings, the degradation is especially pronounced in tasks requiring user patterns understanding, suggesting that such tasks are particularly vulnerable to noise when recalling and executing multi-step procedures grounded in implicit, personalized knowledge.
@@ 5.3.2 Memory Top-K Analysis
Figure 6: Examples of success and error cases in memory utilization stage. Top: success and failure cases of object semantics; Bottom: success and failure cases of user patterns.
5.4 Empirical Analysis of Success and Error Cases
To better understand how LLM-powered embodied agents utilize personalized knowledge derived from episodic memory during task planning, we conduct qualitative case studies of both successful and failed episodes. Figure 6 presents the taxonomy of case types along with illustrative examples. Detailed explanations for each case type are provided in Appendix D.4. Below, we highlight a set of representative cases that offer key insights into the embodied agents’ behavior.
  • Agent misses personalization cues (B.1): The agent fails to recognize user-specific references and treats them as generic or proper nouns, stemming from incorrect interpretations of user intent.
  • Use of commonsense knowledge over personalized knowledge (C.1 & D.1): Even when relevant episodic memory is available, the agent often relies on general commonsense knowledge to infer user routines, rather than using the personalized knowledge. This tendency appears in both successful and failed episodes.
6 Discussion
6.1 The Impact of Trajectory in Episodic Memory
Table 2: Analysis of how memory type affects agent performance: (a) complete action action trajectories containing user instructions, (b) summary of (a), and (c) user instructions only.
@@ 6.1 The Impact of Trajectory in Episodic Memory
GPT-4o (a) 90.0 83.3
(b) 88.0 83.3
(c) 62.4 50.0
Qwen-2.5-72b (a) 77.2 66.7
(b) 77.4 70.0
(c) 51.3 40.0
Llama-3.1-8b (a) 72.8 63.3
(b) 49.4 43.3
(c) 40.0 30.0
Qwen-2.5-7b (a) 50.1 43.3
(b) 43.9 36.7
(c) 35.6 23.3
Memento’s design incorporates complete action-observation trajectories in episodic memory, raising the question of whether agents should reference these detailed trajectories rather than relying solely on user instructions. To investigate this, we conduct an additional comparative experiment to evaluate if providing full action trajectories offers significant advantages over simply providing instructions that state personalized knowledge. For the experiment setup, we evaluate model performance across three different cases of memory provided to the agent as shown in Table 2. The results reveal that while larger models (GPT-4o, Qwen-2.5-72b) perform well with only high-level plans (b), smaller models (Llama-3.1-8b, Qwen-2.5-7b) require full procedural details from completed trajectories (a) to succeed. Most notably, all models show substantial performance drops when given only user instructions (c), suggesting that action trajectories contain essential procedural cues necessary for understanding personalized knowledge, regardless of model capacity. Experiment setup details are provided at Appendix D.5.
6.2 Agent Behavior under Ambiguous Instructions from Users
Model PC (%) SR (%)
GPT-4o (Baseline) 92.0 90.0
GPT-4o 80.4 73.3
Qwen-2.5-72b (Baseline) 75.1 66.7
Qwen-2.5-72b 59.6 53.3
While Memento focuses on evaluating personalized knowledge grounding with explicit references to prior interactions, real-world human-agent communication often involves ambiguous or indirect references. To explore this challenge, we conducted a proof-of-concept experiment to assess whether current models can interpret ambiguous instructions that indirectly refer to previously encountered personalized knowledge.  [^3] Further details of the dataset and experiment setup are provided in Appendix D.5. We created a tailored set of tasks referencing personalized knowledge using contextual cues, synonyms or causal references (e.g., Can you set my afternoon tea time routine? $\rightarrow$ I’m about to enjoy my afternoon tea. Could you set things up as I like them?). The results, shown in Table 3, reveal the degradation in performance, indicating that handling ambiguous queries remains a key challenge for future personalized embodied agents. We view this as a promising direction for future work, where we plan to systematically extend Memento to evaluate LLM-powered embodied agents’ capabilities under ambiguous and implicit reference scenarios, aiming to better reflect the complexity of real-world human-agent interaction.
@@ 7 Conclusion
In this work, we present Memento, a novel personalized embodied agent evaluation framework designed to assess the LLM-powered embodied agents’ ability to utilize episodic memory for providing personalized assistance. Our experiments across a range of LLM-powered embodied agents reveal key limitations in their ability to effectively leverage personalized knowledge from memory, particularly when integrating multiple memories and interpreting user patterns. These findings highlight the gap between current capabilities and the demands of real-world personalized assistance. We hope that Memento serves as a stepping stone for future research in developing more effective personalized embodied agents.
Appendix A Limitations
@@ A.1 Limitations
Controlled simulator environment.
Our experiments are conducted entirely in a controlled simulator environment [38], which does not fully reflect the complexities of real-world robotics such as perception noise, actuation uncertainty, or unstructured environments.
Visual perception is not addressed.
Our framework deliberately isolates memory-centric planning by excluding visual perception components such as VLMs [55] or VLA [6, 24] models. This enables focused evaluation of memory utilization, but limits the system’s generalizability to fully grounded perception scenarios.
Oracle skills.
We intentionally employ oracle skills for low-level perception and motor control to isolate and focus on the high-level planning and memory reasoning capabilities of the LLM-based agents. As a result, our framework does not evaluate the embodied agent’s full task execution process.
A.2 Societal Impacts
@@
Our work explores how embodied agents can remember and adapt to user-specific preferences through episodic memory, enabling more personalized and natural interactions. This capability has the potential to enhance convenience, efficiency, and user satisfaction in everyday environments—benefiting a wide range of users regardless of age or ability. However, personalization also raises concerns about privacy, bias reinforcement, and over-dependence on AI agents [41, 25]. Since episodic memory involves storing interaction history, future systems must consider secure and transparent memory handling. Although our study is conducted in a controlled simulator, these considerations will become crucial as such systems move toward real-world deployment. We hope that Memento serves as a foundation for future research in building safe, privacy-aware, and user-aligned personalized embodied agents.
Appendix B Details of Experiment Setup
B.1 Evaluation Method
We adopt the official evaluation protocol from the PartNR benchmark [7], which provides a python-based framework for assessing multi-step rearrangement tasks. We use this framework without modification. The evaluator analyzes simulator states using three components: (1) propositions that define object relationships to satisfy (e.g., is_on_top([spoon_1], [table_1])), (2) dependencies that define temporal conditions for multi-step instructions (after_satisfied, after_unsatisfied), and (3) constraints that enforce execution requirements (e.g., ordering, object consistency across steps). We rely on this system to evaluate tasks involving ambiguous references and sequential dependencies. The evaluation produces a percent-complete score and binary success indicator.
B.2 LLM-Powered Embodied Agent Architecture

### v2

@@ 3.3 Memento: Personalized Embodied Agent Evaluation Framework
Figure 3: The performance results without using episodic memory. Original indicates episodes from PartNR dataset.
Dataset construction. We construct the dataset for Memento using Habitat 3.0 simulator (Puig et al., 2023) with a simulated Spot robot (Boston Dynamics, 2025; Yokoyama et al., 2023). Our dataset spans 12 scenes with a total of 438 episodes, using PartNR (Chang et al., 2024) test set as foundation data since its multi-object rearrangement tasks align with our evaluation design.
Using GPT-4o, we generate personalized knowledge and create two instruction types: $I_{acq}$ concatenating base instruction, visual captions, and personalized knowledge for direct goal inference; and $I_{util}$ containing personalized references requiring memory recall. To prevent agents from identifying targets without personalized knowledge, we augmented scenes by placing distractors of the same type but with different appearances near target objects (e.g., placing a "red cup" next to the target "blue cup"). For quality control, we filtered episodes with similar memories referencing identical objects and manually reviewed failed cases to ensure instruction quality. We further validate the dataset with an inter-rater reliability analysis, achieving an average Krippendorff’s alpha of 0.69 across four key criteria. Details of the Memento are provided in Appendix C.3.
3.4 Metrics & Retrieval Setup
[…]
Retrieval setup. For memory retrieval in the utilization stage, we use top-$k$ similarity-based retrieval with all-mpnet-base-v2 (Reimers & Gurevych, 2019), treating the current instruction as a query and episodic memory instructions as keys. Since our primary objective is evaluating memory utilization capability rather than retrieval performance, we ensure the corresponding gold memories are always included in the top-$k$ results through randomly replacing retrieved candidates when necessary. We set $k=5$ as our default configuration because this achieves 96.5% recall for gold memories in realistic retrieval scenarios, compared to only 86.1% recall at $k=3$.
4 Evaluating Memory Utilization Across Embodied Agents
Through Memento, we assess memory utilization capabilities across multiple LLM-powered embodied agents of varying parameter sizes to understand how effectively they provide personalized assistance and identify their limitations (RQ1).
4.1 Models & Implementation Details
We evaluate both proprietary models (GPT-4o (Hurst et al., 2024), Claude-3.5-Sonnet (Anthropic, 2024)) and open-source models (Llama-3.1-70b/8b (Grattafiori et al., 2024), Qwen-2.5-72b/7b (Yang et al., 2024)) with different model families and sizes. We implement LLM-powered embodied agents following established architectures (Szot et al., 2021; Puig et al., 2023; Chang et al., 2024), where LLM serves as a high-level policy planner that selects appropriate skills from a predefined skill library using ReAct (Yao et al., 2023) prompting format. Note that, we use gold perception and motor skills to focus on challenges related to embodied agents’ cognitive abilities. Complete implementation details are provided in Appendix C.1.
@@ 4.1 Models & Implementation Details
Model Stage Task Type (Memory) Planning Cycles $\downarrow$ Sim Steps $\downarrow$ Percent Complete $\uparrow$ $\Delta$ PC $\uparrow$ Success Rate $\uparrow$ $\Delta$ SR $\uparrow$
@@
4.2 Results
@@
LLM-powered embodied agents struggle with understanding personalized knowledge. In Table 1, we observe that while GPT-4o maintains a relatively high success rate in the single-memory tasks, all models show a success rate drop over 20% compared to the memory acquisition stage. This substantial performance decline across all evaluated models demonstrates that current embodied agents cannot reliably integrate personalized knowledge for task execution.
Current embodied agents can effectively recall object semantics but struggle to comprehend user patterns. We then analyze the agents performance on single-memory task by personalized knowledge types. Notably, as shown in Figure 4, all models show minimal performance degradation for object semantics tasks, while user pattern tasks exhibit substantial performance drops across all evaluated models, indicating that current embodied agents face significant limitations in comprehending sequential information within behavioral contexts.
Figure 5: Taxonomy of successful and failed cases in memory utilization with illustrative examples. Top: success and failure cases of object semantics; Bottom: success and failure cases of user patterns.
Agents default to commonsense over personalized knowledge. To understand the agents’ behavior on personalized assistance tasks, we conduct qualitative case study which is shown in Figure 5. Through this analysis, we identify an interesting pattern: agents tend to rely on commonsense knowledge rather than personalized memory. For example, when reasoning about "afternoon refresh setup" (C.1), the agent relies on commonsense knowledge rather than referencing personalized memory, correctly predicting "jug and cup" because this commonsense happens to align with the user’s actual behavior. However, when applying the same commonsense-based reasoning approach to "playtime" (D.1), the agent incorrectly places toys on the TV room table instead of consulting the personalized memory that specifies the user’s preference for placing toys on the bench. This indicates that when agents struggle to reference personalized knowledge from memory, they default to parametric commonsense knowledge rather than leveraging non-parametric personalized knowledge, leading to inconsistent performance across user contexts.
5 Scenario-based Analysis for Memory Utilization Bottlenecks
In this section, we conduct scenario-based analysis to understand the factors that negatively affect agents’ memory utilization capabilities in personalized assistance tasks (RQ2). First, we examine agent performance when increasing the number of retrieved memories in the context (§˜5.1). Second, we analyze agent behavior when tasks require coordinating multiple memories to achieve task completion (§˜5.2).
@@ 5 Scenario-based Analysis for Memory Utilization Bottlenecks
5.1 Analyzing Agent Performance Under Increased Memory Context
To understand whether agents remain robust in identifying correct information as memory volume increases, we vary the number of retrieved memories (top-$k$) in single-memory tasks. We ensure gold memory is always included and focus on whether agents can reference the correct memory.
Information overload degrades memory utilization performance. As shown in Figure 6, increasing $k$ consistently degrades performance across all models and both personalized knowledge types, indicating that agents struggle to extract relevant information when presented with larger sets of retrieved memories. We found that agents increasingly rely on commonsense knowledge rather than personalized memories as retrieved memories increase, especially on user patterns, suggesting that embodied agents suffer from information overload. Therefore, even though LLMs demonstrate impressive capabilities in processing long-context information (Wang et al., 2024b; Liu et al., 2025), memory systems that retrieve appropriate numbers of memories are crucial for personalized embodied agents. We provide further analysis of retrieval strategies in scenarios without guaranteed gold memory access in Appendix E.3.
5.2 Investigating Agent Behavior in Joint-memory Tasks
In realistic scenarios, users often refer to multiple pieces of personalized knowledge within a single instruction. Accordingly, we fix the top-$k$ parameter to $k=5$ and employ joint-memory tasks in Memento to analyze the agents’ performance on tasks that require synthesizing information from two distinct episodic memories.
Figure 7: Personalized knowledge type-based analysis on joint-memory tasks.
[…]
6 Exploring Memory Design for Personalized Embodied Agents
Table 2: Model performance with memory format simplification approaches. Full memory indicates the episodic memory.
@@ 6 Exploring Memory Design for Personalized Embodied Agents
GPT-4o Full Memory 90.0 83.3
Summarization 88.0 83.3
Instruction-only 62.4 50.0
Qwen-2.5-72b Full Memory 77.2 66.7
Summarization 77.4 70.0
Instruction-only 51.3 40.0
Llama-3.1-8b Full Memory 72.8 63.3
Summarization 49.4 43.3
Instruction-only 40.0 30.0
Qwen-2.5-7b Full Memory 50.1 43.3
Summarization 43.9 36.7
Instruction-only 35.6 23.3
Based on our previous findings that demonstrate critical bottlenecks in memory utilization for personalized assistance, we identify that these limitations stem from the fundamental challenge of extracting relevant personalized knowledge from episodic memories. We hypothesize that clearly providing personalized knowledge information—an approach consistent with recent advances in context engineering (Mei et al., 2025)—will enhance agents’ ability to leverage memory effectively. Therefore, in this section, we explore memory architectural approaches to improve agents’ memory utilization capabilities for personalized assistance tasks (RQ3).
6.1 Investigating Memory Format Simplification
As episodic memory contains full interaction histories, it includes contextual information irrelevant to personalized knowledge extraction. Therefore, we investigate memory format simplification approaches that provide agents with refined information by removing irrelevant content, hypothesizing that this reduction of noise will improve performance. We compare full episodic memory against two simplified alternatives: (1) GPT-4o summarized episodic memory that retains key information while removing trajectory details, and (2) instruction-only setups that rely solely on task instructions, since personalized knowledge can be extracted from instructions alone.
Episodic memory provides both personalized knowledge and in-context learning benefits. As shown in Table 2, summarized memory shows minimal impact on large models (GPT-4o, Qwen-2.5-72b). Notably, smaller models (Llama-3.1-8b, Qwen-2.5-7b) exhibit performance degradation when using summarized memory compared to full episodic memory. This performance drop occurs because summarizing removes the trajectory information that provides crucial in-context learning benefits for LLMs, reducing the overall task success rate. These findings demonstrate that episodic memory is an essential module for embodied agents.
6.2 Developing Hierarchical Knowledge Graph-based User Profile Memory
Figure 8: Agent performance with user profile memory across single- and joint-memory tasks.
Building upon our finding that episodic memory is an essential memory module, we devise a hierarchical knowledge graph-based user profile memory module that manages personalized knowledge independently to provide cleaner, more accessible personalized information to the agent.
Figure 9: Illustration of user profile memory.
Memory module design. A key challenge in developing a personalized knowledge module is organizing knowledge relationships effectively. Sequential knowledge like user patterns requires sophisticated memory structure to prevent information loss or failing to adapt to evolving knowledge (e.g., "Add the mug from my coffee set to my morning routine and prepare my routine"). Inspired by previous works on knowledge graphs (Speer et al., 2017; Bosselut et al., 2019), we construct personalized knowledge as a hierarchical knowledge graph that captures both object semantics and user patterns.
As shown in Figure 9, user profile memory employs a three-level structure with users at the top, knowledge types (object semantics and user patterns) in the middle, and specific elements (objects, patterns, locations) at the bottom. The graph connects these levels using hierarchical edges for structural relationships and temporal edges for sequential ordering within user patterns. This design enables systematic representation and easy updates of personalized knowledge—for instance, a red mug from a coffee set can be dynamically added to a morning routine by updating the relevant knowledge relationships. More details of the user profile memory design, schema, extraction and retrieval algorithms are provided in Appendix D.1.
User profile memory enhances embodied agents’ memory utilization capability. As shown in Figure 8, agents with user profile memory achieves significant performance improvements across all models for both single-memory and joint-memory tasks. The performance gains are especially notable on user pattern tasks, where the sequential structure had previously posed challenges. These results indicate that user profile memory effectively manages personalized knowledge, demonstrating its ability to enhance agents’ memory utilization capabilities for personalized assistance tasks.
@@ 7 Conclusion
This work investigates memory utilization challenges in LLM-powered embodied agents for personalized assistance through Memento, revealing that while agents can recall simple object semantics, they struggle with sequential user patterns, suffer from information overload when processing multiple retrieved memories, and fail to coordinate joint-memory tasks effectively. We demonstrate that episodic memory provides both personalized knowledge and in-context learning benefits, explaining why simple summarization approaches fail. To address these bottlenecks, we design a hierarchical knowledge graph-based user-profile memory module that separately manages personalized knowledge while preserving episodic memory benefits, achieving substantial improvements across both single-memory and joint-memory tasks. We hope these findings will inspire future research toward more sophisticated memory architectures that enable truly effective personalized embodied agents.
Ethics Statement
This work involves no human subjects, conflicts of interest, or legal compliance concerns. We use licensed resources (e.g., PARTNR (Chang et al., 2024) under the MIT License) and commonsense knowledge without personal data, copyrighted content, or bias-amplifying material. We generated personalized knowledge from LLM-synthesized descriptions with prompts designed to reduce biased or discriminatory content, and we manually verified that the resulting user knowledge was free of such concerns. We acknowledge that embodied memory systems may pose privacy and security risks in real-world deployment. To mitigate these issues, future systems should adopt on-device processing, minimize retention, require user consent, and apply secure, transparent memory handling. We will release only de-identified or synthetic assets and provide documentation of safeguards to reduce misuse risk.
Reproducibility Statement
To facilitate full reproducibility of our results, we include complete evaluation dataset details and experimental procedures in the appendices. Furthermore, we publish our code at https://anonymous.4open.science/r/MEMENTO with thorough documentation.
Appendix A Limitations, Privacy, Societal Impacts and LLM Usage
@@ A.1 Limitations
This work focuses on agents’ memory utilization capabilities for providing personalized assistance. While our study provides valuable insights, several limitations must be acknowledged. First, our experiments were conducted in a controlled simulator environment (Puig et al., 2023). Although we focused on agents’ cognitive abilities and expect marginal impact on our results, this controlled environment may not fully capture the complexity and variability of real-world, presenting potential risks for generalization. Second, we utilize oracle perception and motor skills to isolate memory utilization as the primary performance factor. Recent research has increasingly incorporated visual components into embodied agents (Huang et al., 2023a; Yang et al., 2025) or developed Vision-Language-Action (VLA) models that generate end-to-end actions (Brohan et al., 2023; Kim et al., 2024; Bjorck et al., 2025). Additionally, recent studies have shown that primary limitations in embodied tasks often exist at the observation and action levels (Chang et al., 2024; Sapkota et al., 2025). Therefore, while our focus on language-based memory utilization capabilities provides clear insights into this specific aspect, it may not capture all failure modes present in actual embodied agents. Finally, our framework lacks conversational components that are crucial for effective personalized assistance. Beyond memory utilization, personalized embodied agents require robust communication capabilities, including the ability to clarify ambiguous instructions, proactively engage with users, and adapt their communication style to individual preferences. Investigating how memory utilization integrates with conversational abilities to enable more comprehensive personalized assistance represents an important direction for our ongoing research.
A.2 Privacy and Security Considerations
As our work involves personalized knowledge and agent memory systems, we recognize potential privacy risks including unauthorized access to user-specific knowledge and inference of sensitive personal information from interaction patterns. While our current study utilizes synthetic personalized data to mitigate privacy concerns, real-world deployment of personalized embodied agents requires robust privacy safeguards including encryption of user-specific knowledge bases, secure access controls with proper authentication protocols, data minimization strategies, and mechanisms for user control over personal data. The specific implementation of these privacy protection mechanisms should depend on deployment context and applicable regulatory requirements, and we recommend thorough privacy impact assessments before real-world deployment.
A.3 Societal Impacts
@@
Our work explores how embodied agents can remember and adapt to user-specific preferences through episodic memory, enabling more personalized and natural interactions. This capability has the potential to enhance convenience, efficiency, and user satisfaction in everyday environments—benefiting a wide range of users regardless of age or ability. However, personalization also raises concerns about privacy, bias reinforcement, and over-dependence on AI agents (Rueben & Smart, 2016; Lee et al., 2012). Since episodic memory involves storing interaction history, future systems must consider secure and transparent memory handling. Although our study is conducted in a controlled simulator, these considerations will become crucial as such systems move toward real-world deployment. We hope that Memento serves as a foundation for future research in building safe, privacy-aware, and user-aligned personalized embodied agents.
A.4 LLM Usage
Our use of LLMs was for the following specific purposes:
  • Writing and Editing Assistance: We utilize LLMs for improving the clarity, grammar, and readability of the text. This included correcting grammatical errors, and ensuring consistent terminology throughout the paper.
  • Code Generation for Experiments: We utilize LLMs to accelerate the implementation of our experimental framework.
The authors meticulously reviewed, edited, and validated all content generated or modified by LLMs. The conceptual arguments, experimental design, analysis, and final conclusions presented in this paper are the original work of the authors. We take full responsibility for all content of this work.
Appendix B Extended Related Work
Our work focuses on memory and personalization in embodied agents. Here, we broaden the related work to include studies on memory-augmented agents, personalization, and memory architectures within the wider research context.
Memory-augmented agents have increasingly been studied as a means to support long-term reasoning, planning, and adaptive behavior in LLM agents. Park et al. (2023) propose Generative Agents, which simulate human-like behavior by maintaining a memory stream of past experiences in natural language. This memory stream enables agents to reflect on prior events, retrieve relevant information, and plan future actions based on their histories. In a similar direction, Xu et al. (2025) introduce A-Mem, a dynamic memory system in which memory is represented as evolving and interconnected notes. The agent can generate, retrieve, and update these notes over time, thereby supporting adaptability.
Memory for dialogue and personalization has emerged as a key direction, highlighting how memory supports user-specific context and long-term consistency in interactions. Li et al. (2025a) present a system that leverages both short-term and long-term memories to maintain user-specific context across multiple conversation sessions, reporting improvements in dialogue coherence and persona consistency. Meanwhile, Wang et al. (2024a) propose the Self-Controlled Memory (SCM) framework, where the agent dynamically decides when and what to store or retrieve, achieving greater coherence and knowledge retention over extended interactions. Alongside these approaches, structured long-term and episodic memory have been introduced to enhance personalization. Zhong et al. (2023) develop MemoryBank, which stores and retrieves interaction history to support consistent responses across turns. Das et al. (2024) present Larimar, a framework that enables selective recall and forgetting, providing controllable knowledge updates that can support personalization at scale.
Architectural perspectives on memory focus not only on building memory components for individual tasks but on treating memory as a fundamental part of the overall system. Packer et al. (2024) illustrate this idea with MemGPT, which frames memory management through an operating-system analogy. In this view, the LLM functions like an operating system that autonomously organizes and manages both its in-context and external memory, rather than relying solely on user-provided prompts. This perspective emphasizes memory as a central architectural layer, enabling scalable and autonomous agents that can handle complex, long-horizon tasks.
Appendix C Details of Evaluation Setup
C.1 LLM-Powered Embodied Agent Architecture